6. Consider the following balanced equation MnO2 + 4H+ + 2Cl- → Mn2+ + 2H20 +Cl2

4.786g MnO2 are reacted and the chlorine gas is collected in a 458 mL vessel at 19°C. What is the pressure of the gas?
Answer #1

4.786 g MnO2 = 4.786 / 86.94 g/mol = 0.055 mol MnO2

As per the equation 1 mol of MnO2 gives 1 mol Cl2 gas

0.055 mol MnO2 gives = 0.055 mol Cl2 gas = n

V = 458 mL = 0.458 L

R = 8.314

T = 19 +273K = 292 K

PV = nRT

P = nRT / V = 0.055 *0.08206 * 292 / 0.458 = 2.88 atm